OpenStudy (anonymous):

Algerbra 1. p.129 #17 PHIYSIS accceleration is the measure of how fast a velocity is changing. The formula foracceleration is a=vf-vi over t. a represents the acceleration rate, vf is the final velocity, vi is the intial velocity, and t represents the time in seconds. a. solve the fomula for vf. i got a*t+vi=vf b. What is the final velocity of a runner who is accelerating at 2 feel per second squared for three seconds with an initial velocity of 4 feet per second? What do i do and is the top answer right!??

OpenStudy (anonymous):

\[v _{f}=4+2(3)\]

OpenStudy (anonymous):

@Chlorophyll can u help me with this? or anyone else who is willlign to help (:

OpenStudy (anonymous):

convert answer from feet to metres

OpenStudy (anonymous):

@Jonask how did u get there??!

OpenStudy (anonymous):

\[v _{f}=v _{i}+at\]

OpenStudy (anonymous):

\[v _{i}=4,a=2,t=3\]

OpenStudy (anonymous):

a. solve the fomula for vf. a = ( vf -vi ) t. ->at = vf -vi =>vf = at + vi

OpenStudy (anonymous):

so you substitute this values

OpenStudy (anonymous):

@smile.im.here. clear enough?

OpenStudy (anonymous):

@Chlorophyll that's bacily wat i got for a. but what am i suposed to do for b.?

OpenStudy (anonymous):

b. What is the final velocity, vf : Given a = 2 ft/sec² t =3 secs, vi = 4 ft/sec

OpenStudy (anonymous):

=>vf = vi + a t All you do is to plug the numbers in to get vf.

OpenStudy (anonymous):

yeah..i like algerbra and all but i cant read it. so this is just a gues..

OpenStudy (anonymous):

\[v _{_{f}}=4+2^{3}\] wud that be right?

OpenStudy (anonymous):

and @Chlorophyll yes i meant 'guess'

OpenStudy (anonymous):

what? "can't read"?

OpenStudy (anonymous):

OHHH!! yeah! no, i dont really understand it if im not in the room with the teacher. heh..

OpenStudy (anonymous):

Now just match the numbers with the letters properly: vf = vi + a* t vi = 4 ft/sec a = 2 ft/sec² t =3 secs,

OpenStudy (anonymous):

so wud the equation i put up be correct? and then just solve for the final velocity, right?

OpenStudy (anonymous):

YAY!

OpenStudy (anonymous):

tks (:

OpenStudy (anonymous):

just a second.

OpenStudy (anonymous):

do i combine 4ft/sec and 2ft/sec? and then divide by three?

OpenStudy (anonymous):

vf = vi + a* t ( vi = 4 ft/sec, a = 2 ft/sec², t =3 secs) = 4 + 2 * 3 There's no divide at all!

OpenStudy (anonymous):

but...ohh, i think my brain wants to move a variable. ok 4+2=6*3=18. i think. im not very good at this stuff

OpenStudy (anonymous):

vf = vi + a* t ( vi = 4 ft/sec, a = 2 ft/sec², t =3 secs) = 4 + ( 2 * 3 ) Always multiply first!

OpenStudy (anonymous):

oh so then vf=10

OpenStudy (anonymous):

i told u i was not great at this stufff (:

OpenStudy (anonymous):

Good luck with your algebra study, then :)

OpenStudy (anonymous):

oh yess i know and we have eca's this year. fun

OpenStudy (anonymous):

and thanks for helping :D